OpenStudy (anonymous):

Factor the expression below.81x^2-100

OpenStudy (cwrw238):

difference of 2 squares 9x and 10 are sorts

OpenStudy (cwrw238):

* square roots

OpenStudy (anonymous):

To factor 81x2 - 100 use difference of squares formula, because both terms are perfect squares. The formula is a2 - b2 = (a - b)(a + b) In this example we have

OpenStudy (anonymous):

81x2 - 100 = (9 x)2 - 102 = (9 x - 10)(9 x + 10)

OpenStudy (anonymous):

thank you so much